when 50% of a number is subtracted from 70, the result is 2 less than the number
write equation and solve

Answer by stanbon(75887)

when 50% of a number is subtracted from 70, the result is 2 less than the number
write equation and solve
-----------
Let the number be "x":
Equation:
70-0.5x = x-2
(3/2)x = 72
x = (2/3)72
x = 48
===============
Cheers,
Stan H.